OpenStudy (anonymous):

lim x approaches pi (1-sin(x/2)/pi-x) cant use LHospitals rule ;/

hartnn (hartnn):

tried substituting y = pi-x ? as x -> pi x-pi ->0 y->0

OpenStudy (anonymous):

0/0 ..

OpenStudy (anonymous):

come on noone knows this?

hartnn (hartnn):

what did you get for sin(x/2)in terms of y ?

OpenStudy (anonymous):

sin(pi-y/2) ??if x=pi-y

hartnn (hartnn):

you mean sin ((pi-y)/2) ? which is sin (pi/2 - y/2) what does that simplify to ?

OpenStudy (anonymous):

whaat? sorry I am not that good in english? do you, by simplify, mean that there is some kind of trig equation that is egual to sin(pi/2-y/2) ?? because if you ment to input y->0 the result is again 0/0

hartnn (hartnn):

yes, i mean there is a simplified trigonometric expression for sin (pi/2-y/2)

hartnn (hartnn):

sin (pi/2 - theta) = cos theta

OpenStudy (anonymous):

ok let me try thanks

OpenStudy (anonymous):

nope still 0/0

OpenStudy (anonymous):

Something that might also be useful: \(\cos x=-\cos(x-\pi)\).

hartnn (hartnn):

don't plug in 0 now... unless you cancel out a common factor from numerator and denominator (or using standard formulas like sin x/x =1)

hartnn (hartnn):

what limit you have till now ?

hartnn (hartnn):

lim y->0 cos (y/2)/y right ?

hartnn (hartnn):

[1-cos(y/2)]/y

OpenStudy (anonymous):

tried all that and every supstitution there is..working on this for two days... right

hartnn (hartnn):

ok then multiply and divide by 1+cos(y/2) and try to simplify the numerator (1-cos(y/2))(1+cos(y/2))

OpenStudy (anonymous):

Try it a slightly different way (with same result): \[\large\begin{align*}\lim_{x\to\pi}\frac{1-\sin\dfrac{x}{2}}{\pi-x}&=\lim_{x\to\pi}\frac{1-\sin\dfrac{x}{2}}{\pi-x}\cdot\frac{1+\sin\dfrac{x}{2}}{1+\sin\dfrac{x}{2}}\\\\ &=\lim_{x\to\pi}\frac{1-\sin^2\dfrac{x}{2}}{(\pi-x)\left(1+\sin\dfrac{x}{2}\right)}\\\\ &=\lim_{x\to\pi}\frac{1-\dfrac{1-\cos x}{2}}{(\pi-x)\left(1+\sin\dfrac{x}{2}\right)}\\\\ &=\frac{1}{2}\lim_{x\to\pi}\frac{1+\cos x}{(\pi-x)\left(1+\sin\dfrac{x}{2}\right)}\\\\ &=\frac{1}{4}\lim_{x\to\pi}\frac{1+\cos x}{\pi-x}\\\\ &=\frac{1}{4}\lim_{x\to\pi}\frac{1-\cos (x-\pi)}{\pi-x}\\\\ &=\frac{1}{4}\lim_{x\to\pi}\frac{1-\cos (\pi-x)}{\pi-x}\end{align*}\] which strongly resembles a known limit, \[\large\lim_{x\to c}\frac{1-\cos (x-c)}{x-c}=\color{red}{?}\]

OpenStudy (anonymous):

0/0 again??

OpenStudy (anonymous):

... There's no way to tell you what the next step is without telling you the answer outright.

OpenStudy (anonymous):

\[\text{Do you know what }\lim_{x\to0}\frac{\sin x}{x}\text{ is?}\]

OpenStudy (anonymous):

Hint: NOT 0/0

OpenStudy (anonymous):

i know that the answer is zero and i know to get to it by LHospitals rule but i do not know how to get a zero without the rule

OpenStudy (anonymous):

And you don't need L'Hopital's rule. You just need to recall a *known* limit and be able to apply it here.

OpenStudy (anonymous):

in line 5, how did you get 1/2 out of (1+sin (x/2)) ?

OpenStudy (anonymous):

i mean 2

OpenStudy (anonymous):

I used the property that\[\lim_{x\to c}f(x)g(x)=\lim_{x\to c}f(x)\times\lim_{x\to c}g(x)\]

OpenStudy (freckles):

you could also recall the definition of derivative \[\lim_{x \rightarrow \pi}\frac{f(x)-f(\pi)}{x-\pi}=f'(x)|_{x=\pi} \\\text{ and you have something that looks similar } \\ \lim_{x \rightarrow \pi}\frac{\sin(\frac{x}{2})-1}{x-\pi}\]

OpenStudy (anonymous):

cant figure this out.... ://study electrical engineering they said, it will be fun they said...

OpenStudy (anonymous):

Which method do you prefer? Using algebra to rewrite the limit so that it simplifies to something you (should) know, or the definition of the derivative as per freckles's excellent suggestion? (The second is much more elegant in retrospect)

OpenStudy (anonymous):

the first one please :D

OpenStudy (anonymous):

love algebra :D

OpenStudy (anonymous):

Alright, well if you can follow up to the last step, then you have a solid understanding of the algebraic manipulation involved here. The very last step is calculus, or rather remembering things you would usually learn in a calc course.

OpenStudy (anonymous):

we dont have that kind of subjects on our university only mathematichs which implies everything tigether, algebra, geometry, functions and so on..but I will try to find solution to final step.. thanks guys :)

OpenStudy (anonymous):

You don't get offered a course in calculus, yet you have an engineering department? Suspicious.